OpenStudy (anonymous):

could this be 2a-5b? Simplify the following expression. a + c − 4b − c + a − b

OpenStudy (marissalovescats):

Well combine like terms!

OpenStudy (marissalovescats):

a+a=? +c-c=? -4b-b=? You are correct! 2a-5b!

OpenStudy (anonymous):

@marissalovescats thank you for the help(: